OpenStudy (amtran_bus):

.444444=.666667^x Solve for x.

OpenStudy (amtran_bus):

I know it is 2 but wolfram makes you pay to see steps now!

OpenStudy (anonymous):

Ya ik they implemented that a year ago :(

OpenStudy (anonymous):

Are you familiar with log

OpenStudy (amtran_bus):

Yes. Would I take the log of both sides?

OpenStudy (anonymous):

Yes \[\log(.444444)=\log(.666667^x)\] \\[log(.444444)=x\log(.666667)\] \[\frac{\log (.444444)}{\log (.666667)}=x\]

OpenStudy (amtran_bus):

Thank you so much!